A tsunami wave travels at a sped of 159 km/hr, the amplitude of the wave inside the ocean is 200 m and the distance from crest to crest is 48 km. calculate the energy carried by the tsunami


1
Expert's answer
2021-09-12T19:01:38-0400

The energy of gravity waves is given as follows (see https://farside.ph.utexas.edu/teaching/336L/Fluid/node150.html):


"E = \dfrac{1}{2}\rho ga^2"

where "\rho = 1020kg/m^3" is the density of sea water, "g\approx 9.81m/s^2" is the gravitational acceleration, and "a = 200m" is the amplitude of the wave. Thus, obtain:


"E = 0.5\cdot 1020\cdot 9.81\cdot 200^2 \approx 10^{6}J"

Answer. "10^{6}J".